OpenStudy (anonymous):

Isis bought seven boxes. A week later half of all her boxes were destroyed in a fire. There are now only 22 boxes left. With how many did she start?

OpenStudy (pink.1d):

Assume she had x boxes to start with. After buying 7 she had x+7 boxes She lost (1/2)(x+7) in the fire. --- Equation: x - (1/2)(x+7) = 22 x - (1/2)x - (7/2) = 22 (1/2)x - (7/2) = 22 (1/2)x = 51/2 x = 51 boxes

OpenStudy (anonymous):

Three basketball teams measured the height of each player on their team. The table shows the mean and the mean absolute deviation of the heights for each team. Which statements are correct? Choose exactly two answers that are correct. Team Mean height (cm) MAD (cm) Bulldogs 165 6.9 Panthers 177 7.1 Warriors 176 4.8 A. The heights of the Bulldogs’ players vary more than do the Warriors’ heights. B. The heights of the Warriors’ players vary more than do the Panthers’ heights. C. The heights of the Bulldogs’ players and those of the Panthers vary about the same amount. D. The heights of the Panthers’ players and those of the Warriors vary about the same amount.

OpenStudy (pink.1d):

@mayah_abdul15 post that as your own question not on eomone elese's question

OpenStudy (pink.1d):

@xisaaaiahx yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

She Can post on here

OpenStudy (pink.1d):

@xisaaaiahx yes but more people can see it if it is its own question :)

OpenStudy (pink.1d):

@xisaaaiahx can u give me a fan and medal for my answer?